what is the length of one side, square whose area is 81 cm
please tell me the formula for the answer

Found 2 solutions by edjones, stanbon:
Answer by edjones(8007)

sqrt(81)=9 cm

Answer by stanbon(75887)

What is the length of one side of a square whose area is 81 cm^2.
Please tell me the formula for the answer.
-----
Formula: Area of a square = (side)^2
A = s^2
---
Equation:
81 = s^2
s = 9 cm
==================
Cheers,
Stan H.